# LOLA: LLM-Assisted Online Learning Algorithm for Content Experiments

### v1

@@ 3.2.2 Implementation Details
In our numerical study, we employ the embedding model from OpenAI to facilitate a fair comparison with GPT-based prompts. At the time of this study, the latest and best-performing embedding model available was text-embedding-3-large. This model generates embeddings with up to 3,072 dimensions and supports a maximum token limit of 8,191, with its knowledge base extending up to September 2021. A notable feature of this model is its flexibility, allowing users to tailor the embedding dimension to their specific needs. [^11] Other models capable of text embedding, like BERT (devlin2018bert) and Sentence-BERT (reimers2019sentence), have fixed embedding vector dimensions that remain constant once the model is trained. These dimensions cannot be changed or resized for different applications. However, OpenAI’s embedding model uses a technique that, while producing outputs of fixed length, encodes information at different granularities (kusupati2022matryoshka). This allows a single embedding to adapt to the computational constraints of downstream tasks. More specifically, downstream tasks can use only the first several dimensions of the original vector, and the information retained decreases smoothly as a shorter part of the vector is used. In our experiments, we consider both 3,072- and 256-dimensional embeddings to examine whether the size of the embedding vector size affects the performance of the downstream classification models. [^12] A higher dimension does not necessarily mean better performance. While a higher dimension can retain more information for each sample, it also means that the samples are spread more sparsely in a higher-dimensional space given a limited number of training samples. This sparsity can increase the possibility of over-fitting.
Figure 3 illustrates the conceptual framework of our approach here. First, we call the OpenAI embedding API in Python to obtain text embeddings for the headlines based on text-embedding-3-large. We then concatenate the embeddings of two headline embeddings and apply the downstream classification models to predict the probability that the first headline will generate a higher CTR. To assess the impact of model complexity on performance, we test two classification models: Logit and MLP. The embedding model parameters remain fixed during training, while only the classification model parameters are updated.
Figure 3: The pipeline of the classification with the LLM-based text embeddings. Headline 1 and headline 2 are natural language sentences while Embedding 1 and Embedding 2 are numerical vectors.
As discussed earlier, we chose all headline pairs that are significant at 0.05 level for our analysis. It is worth noting that the same headline can appear in multiple pairs when there are more than two headlines tested in one experiment. To avoid information leakage, we first randomly split the data into 80% training and 20% test sets based on the test ID (the experiment ID for Upworthy) to ensure all headline pairs under the same experiment are either in the training set or the test set. Notice that Upworthy allowed additional experiments for the same headline under a new experiment ID because the editor may require an extra test to make an informed decision. It means that we may have the same headlines in both the training and test sets, even if they were in totally different experiments. In such cases, we discard headlines that appeared in the training set from the test set to avoid information leakage. [^13] There are alternative ways to do the train-test split without discarding any samples. One approach is assigning one headline pair at one time to training or test groups and then adding all pairs with the same headline in the same group until all samples are allocated. However, in our study, after conducting several experiments, we found that the splitting procedure has minimal impact on the final results. Hence, to maintain clarity and simplicity in our exposition, we simply discard duplicated testing samples, which constitute around 25% of the original test set.
One may still be concerned if this Upworthy data has been used as part of OpenAI’s training corpus, which could potentially invalidate our analysis. However, we argue this is unlikely for several reasons. First, it is improbable that OpenAI would have specifically downloaded this Upworthy csv format dataset from a subfolder in Upworthy Research Archive. Second, the original data only contains impressions and clicks for each headline, as shown in Table 1, rather than paired headlines with winner information. Therefore, during the training of OpenAI models, it is unlikely they were exposed to the winner information since the training objective is to predict the next word. Third, we conducted an experiment to determine if this dataset was indeed a part of OpenAI’s corpus. We used embeddings from two headlines with significantly different impressions (possibly from different tests) for the downstream task of predicting which headline received higher impressions. If the embedding model was trained on this corpus, its embeddings should reflect the number of impressions and perform well on this task. However, the observed accuracy was approximately 50%, indicating that the embeddings do not contain such information, confirming our assumption.
@@
Finally, as discussed earlier, we consider both a Logit model (with linear features) and a MLP for the classification task (i.e., to predict which headline is catchier). The MLP model employs a fully connected neural network with ReLU activation for hidden layers and Sigmoid transformation for the output probability, with dropout layers added to prevent over-fitting. The MLP network is trained using the stochastic gradient descent algorithm to minimize binary cross-entropy. We perform hyper-parameter tuning with Bayesian Optimization to determine the optimal configuration. For the OpenAI-256E model, we used the following MLP architecture: Inputs(512)-fc(256)-ReLU-dropout(0.5)-fc(1)-sigmoid, and the training hyperparameters were set with an $L_{2}$ penalty of $10^{-4}$ and 74 training epochs. For the OpenAI-3072E model, the MLP architecture was Inputs(6144)-fc(2048)-ReLU-dropout(0.5)-fc(1)-sigmoid, with an $L_{2}$ penalty of $10^{-4}$ and 32 training epochs. Please see Web Appendix A.2 for additional details of this model.
@@ 3.2.3 Performance Analysis of Embedding-based Approaches
We now present the performance of the different embedding-based classification models in Table 6. We also include a benchmark using logistic regression with Word2Vec word embeddings for a comprehensive comparison. We use Word2Vec as it was one of the best embedding models before the wide adoption of Transformer in NLP. Unlike LLM-based text embeddings, Word2Vec does not use Transformers and cannot pay attention to surrounding token information. We compute the Word2Vec embedding vector for each word in the headline and average them to get the headline embedding, a common practice in Word2Vec. To ensure consistency with the dimensions used for OpenAI embeddings, we consider Word2Vec embeddings with 256 and 3,072 dimensions.
@@
OpenAI-256E Logit 77.56%
MLP 78.29%
OpenAI-3072E Logit 82.33%
MLP 82.06%
Word2Vec-256E Logit 68.58%
Word2Vec-3072E Logit 68.84%
@@
Table 6: The performance of different embedding and classification models on test data. All models use the same training data, and accuracy is evaluated on the test data. Note that Logit model’s training is a deterministic process after fixing the data, while MLP training incurs randomness via stochastic gradient descents. Therefore, we rerun MLP training three times with different sample paths (training set shuffling) and report the average of the accuracy from three reruns: 78.27%, 78.29% and 78.32% for OpenAI-256E + MLP and 82.43%, 81.67% and 82.09% for OpenAI-3072E + MLP.
@@
We find that classification models using the 3072-dimensional text embedding from OpenAI achieve more than 80% accuracy, demonstrating superior performance compared to prompt-based approaches. This performance can be attributed to the fact that embedding-based classification models are trained on a customized dataset, allowing it to learn decision rules tailored precisely fine-tuned to the headline selection task. Because these models can leverage a larger labeled dataset during training, they are more generalizable and robust. In contrast, in-context learning relies on the LLM’s ability to understand and generalize from only a few demonstrations, resulting in sub-optimal performance compared to a task-specific trained classifier. Furthermore, the task of choosing the catchier headline from two options may not be a common scenario in OpenAI’s training corpus. Consequently, GPT models may struggle to generalize effectively from the limited examples provided in prompts, leading to less accurate predictions.
Next, we discuss the detailed comparison between models using the embedding technique. We first compare the results under different embedding dimensions. Typically, more dimensions give greater quality encoding and, thus, higher prediction accuracy. However, there is some limit beyond which there are diminishing returns. In our experiment, the 3072-dimensional embedding (OpenAI-3072E) improves over the 256-dimensional embedding (OpenAI-256E) by around 4.3%. It is notable that OpenAI embeddings with Logit outperform Word2Vec significantly by around 9.0% and 13.5% at the same dimensions of 256 and 3072, respectively. These observations are unsurprising since OpenAI’s LLMs are state-of-the-art language models that naturally perform better than the classic Word2Vec embeddings.
A more striking observation is that classification model complexity does not improve the prediction (conditional on the same inputs). Specifically, there is no significant accuracy difference between MLP and Logit when using OpenAI-256E or OpenAI-3072E. This observation suggests that a linear classifier on top of fixed LLM embeddings could effectively capture the attractiveness of content. An interesting question is why text embeddings obtained from models trained for the next word/token prediction task with the cross-entropy loss are able to produce informative features for our downstream classification task. (saunshi2020mathematical) provide an explanation by demonstrating that classification tasks can be reformulated as sentence completion tasks, which can be further solved linearly using the conditional distribution over words following an input text. Thus, the embeddings generated from a next-word prediction task inherently capture the contextual information and relationships between words, making them useful for various downstream tasks such as sentiment analysis and headline selection (as our experiments illustrate).
We further investigate two specific aspects of embedding-based approaches. First, we examine the conditions under which these approaches fail, and second, we quantify the role of the number of training samples on predictive accuracy. For both these analyses, we use logistic regression with linear features as our classification model (since this is the model with the highest accuracy and there is no significant improvement in the accuracy of MLP).
To address the first question, we first calculate the cosine similarity based on the text embeddings for each pair of headlines. This metric captures the similarity between two vectors in inner product space and is defined as the dot product of the vectors divided by the product of their lengths, i.e., ${e_{1}\cdot e_{2}}/(\|e_{1}\|\|e_{2}\|)$, where $e_{1}\cdot e_{2}$ is the dot product of the embedding vectors, and $\|e_{1}\|$ and $\|e_{2}\|$ are the their Euclidean norms. Figure 5(a) shows the distribution of cosine similarity across all headline pairs, whereas Figure 5(b) shows the distributions of cosine similarity separately for two cases: (1) cases where our embedding-based classification correctly predicts which headline is catchier, and (2) cases where our embedding-based classification incorrectly predicts which headline is catchier. We observe a rightward shift in the distribution of incorrect predictions (compared to the distribution of correct predictions). This suggests that incorrect predictions are more common when the headlines are very similar to each other (i.e., higher cosine similarity). There are two possible reasons for this. First, in cases where the headline pairs are very similar, there is a higher chance of Type 1 error (in the t-tests based on our A/B test data), i.e., the headlines may not be significantly different in fact. [^14] Conducting multiple tests in our data can inflate the overall Type I error rate, known as the multiple testing problem. This means there is a higher likelihood of encountering at least one Type I error. Similar headlines are more likely to lead to these errors because the actual difference in their CTRs is likely to be small. This increases the probability that variation in randomly sampled impressions and clicks in A/B tests might lead to significant results by chance, even when there is no true effect. Second, text embeddings may not be able to capture the differences between two headlines when they are similar, i.e., the embeddings may also be very similar. As a result, it is hard for the classification model to differentiate between them. [^15] Note that it is unlikely that this phenomenon is due to problems/issues with the classification model because both Logit and MLP achieve nearly identical performance.
(a) Performance of OpenAI-256E
(b) Performance of OpenAI-3072E
Figure 4: Performance of Logit classification models with OpenAI embeddings under varying training dataset sizes and different embedding sizes.
Next, to address the second question, we conduct experiments varying the proportion of data used for training vs. test. We use the Logit model for these experiments since it offers the best performance and is easy to train. The results from this exercise are shown in Figure 4. As we can see, when we use the 256-dimensional embeddings, approximately 65% of the training set, comprising approximately 25,000 pairs (with the remaining 35% used for the test set), is adequate to achieve high performance. However, when we use the 3072E-dimensional embedding, there is a clear over-fitting issue even if we use the simplest Logit model and 100% of the data for training. This suggests that for larger embedding vectors, we need much larger training datasets. [^16] It is unlikely that this is due to issues with the classification model since the Logit model is inherently parsimonious. To further investigate this issue, we experimented with both $L_{1}$ and $L_{2}$ regularization to control over-fitting. [^17] This was done using the penalty and C control parameters in the LogisticRegression function from sklearn package. Unfortunately, regularization did not improve the accuracy on the test data, suggesting that over-fitting is unlikely to be due to learning noise in the training data. Recall that the average headline in the Upworthy tests received 3,569 impressions and 54.26 clicks. Our analysis suggests that we need much larger A/B tests to learn the relationship between a headline’s text embeddings and its attractiveness effectively. In summary, these comparisons suggest that data samples from Upworthy are insufficient to unlock the potential of LLMs fully. Nevertheless, diverting more and more traffic to A/B tests is costly, and this observation is one of the motivations for our approach later in Section 4.
(b) Cosine similarity over incorrect/correct predicted test pairs
Figure 5: The left panel shows the distribution of cosine similarity for all pairs. The right panel shows the distributions of cosine similarity for both incorrect and correct predictions using the logit classification model with OpenAI-3072E embeddings.
Finally, we discuss the time and compute costs of embedding-based approaches, and contrast them with prompt-based approaches. We find that obtaining text embeddings through the OpenAI API is extremely efficient in terms of time usage and is also more economical compared to direct prompt-based interactions. When using the embedding API, data transmission involves smaller, fixed-size vectors instead of variable-length text, which reduces data overhead and accelerates processing speed. We spent less than \$1 to obtain the text embedding for all 77,245 headlines, while the prompt experiment with only 10,000 pairs cost us around \$20.181818At the time of this study (May 2024), GPT-4 (turbo version) prompt costs \$10 per 1M input tokens and \$30 per 1M output tokens. Further, the embeddings used in our study, text-embedding-3-large, only cost \$0.13 per 1M input tokens and there is no output token cost associated with embeddings. Our entire dataset comprises 77,245 headlines, with an average of 19 tokens per headline, resulting in a total embedding cost of approximately \$0.18. In addition, we found that it is much faster to process embeddings (more than 3,000 pairs per minute), while the prompt experiment is much slower (can only test around 83 pairs per minute). This efficiency can translate into lower operational costs and help with easier integration into real-time applications. Another key advantage of the embedding-based approach is its transparency and robustness. Using simple models like logistic regression with text embeddings from OpenAI can give researchers and managers insight into the training process and source of performance improvements. This transparency is crucial in business settings, where a clear understanding of a model’s behavior and performance is often necessary to mitigate potential risks and biases. In contrast, in-context learning is a black-box approach where the user/manager cannot obtain any insight into the source of the performance improvement.
3.3 Fine-Tuning Open-Sourced LLMs with LoRA
@@
In Section 3.2, we trained classification models using text embeddings as fixed inputs, bypassing the need to train the LLM itself. The performance of these embedding-based approaches reached 82% accuracy without modifying any parameters in the LLM. This raises a natural question: if we can train LLMs directly, can we achieve better performance? However, training LLMs from scratch is costly and resource-intensive. Therefore, in this section, we turn to fine-tuning techniques, which involve partially adjusting the parameters in the LLMs to improve performance while keeping the computational cost manageable. Note that to fine-tune LLMs, we need to be able to access the underlying model parameters. This is only feasible for open-source LLMs, like BERT and LLama, and not for proprietary LLMs like OpenAI’s GPT and Google’s Gemini. The goal of fine-tuning is to adapt a general-purpose LLM for specific downstream tasks through a process typically involving supervised learning using labeled datasets.
@@
For the fine-tuning task, we use Meta’s Llama-3 (meta_llama_3) as our primary base model because it is regarded as one of the most advanced open-source LLMs to date (llama3_model_card). zhao_etal_2024 consider a large-scale experiment where they fine-tune different open-source LLMs and show that for NLP tasks, including news headline generation, Llama-3 offers one of the best performance. [^19] Before the widespread adoption of the term “LLM”, BERT was considered among the best language models available, and is still widely used as a benchmark. To provide a comparison with these earlier models, we also fine-tuned DistilBERT (sanh2019distilbert)—a distilled version of BERT that retains over 95% of BERT’s performance while containing 40% fewer parameters (devlin2018bert). Details on the implementation and performance of the fine-tuned DistilBERT are shown in Web Appendix A.3. Llama-3 offers two versions: one with 8 billion parameters (Llama-3-8b) and another with 70 billion parameters (Llama-3-70b). We opted for the smaller version since it is easier to fine-tune and, as we will see, even with this smaller model, we see evidence of over-fitting (see Section 3.3.4 for details). Hence, we do not consider fine-tuning the 70-billion-parameter model, where over-fitting problems are likely to be exacerbated. Our numerical experiment shows LoRA fine-tuned Llama-3-8b achieves nearly 83% accuracy.
@@
The rest of this section is organized as follows. We first introduce the Transformer block, the fundamental component of most LLMs, including Llama-3, and the basic architecture of Llama-3 in Section 3.3.1. Next, we present the LoRA method for efficient fine-tuning in Section 3.3.2. The goal is not to delve into every technical detail but to provide the necessary background information for a high-level understanding of fine-tuning. Finally, we detail our implementation in Section 3.3.3 and report our results in Section 3.3.4.
3.3.1 Transformer and Llama Architecture
The Transformer model, introduced by (vaswani2017attention), has revolutionized the field of NLP with its innovative attention mechanism and highly parallelizable architecture. Unlike traditional recurrent neural network (RNN) based models, the Transformer uses self-attention mechanisms to process input data. This shift allows for greater computational efficiency and significantly improved performance across a range of NLP tasks.
As illustrated in the right portion of Figure 6, the Transformer block comprises multiple layers, each featuring two primary components: a multi-head self-attention mechanism and a position-wise fully connected feed-forward network. Key enhancements include layer normalization (ba2016layer) and the use of residual connections (he2016deep), which improve training stability and help prevent issues such as vanishing gradients. While some specifics are omitted to maintain brevity and readability, it’s important to note that such architectures are subject to rapid evolution and vary across different LLMs.
Figure 6: Illustration of the Transformer block and architecture typical in most LLMs.
The right portion provides a detailed view of a single Transformer block, internal components included. We omit the depiction of multi-head attention for brevity. Note that different LLMs may use slightly different designs.
Self-Attention Mechanism: The most important innovation in the Transformer block is the self-attention mechanism. It enables each token in a sentence to “pay attention” to all other words when computing its own representation. This is achieved using an attention mask that guides which tokens should influence each other, thereby allowing the model to capture dependencies between words regardless of their distance in the text. Specifically, for a given input sentence, the model first converts each token into an embedding, [^20] This embedding differs from the text embedding used in the previous section. Each embedding here simply corresponds to one token without paying attention to the whole input sequence. The text embeddings used in the previous section are essentially the contextualized embeddings, which are generated by passing the tokens through multiple layers of Transformers, capturing rich contextual information and summarizing the whole input sequence. forming a matrix $X$ where each row corresponds to the embedding of a token. If the sentence contains $n$ tokens, and each token is represented by a $d$-dimensional embedding, then $X$ is an $n\times d$ matrix. From these embeddings, the model generates three matrices: Query $Q$, Key $K$, and Value $V$:
$Q=XW_{Q},\quad K=XW_{K},\quad V=XW_{V}$
where each head computes an independent self-attention process:
$\text{head}_{i}=\text{Attention}(XW_{Q}^{i},XW_{K}^{i},XW_{V}^{i}),$
$\text{FFN}(x)=\max(0,xW_{1}+b_{1})W_{2}+b_{2}$
It is worth noting that LLMs do not necessarily operate in a conversational manner, nor do they always use language as their output. As we will demonstrate in Section 3.3.3, we employ the Llama model as a classifier, where the output is the distribution of predicted classes. In applications like conversational AI, exemplified by GPT, the contextualized embeddings from the last token of the last layer of the Transformer are fed into a linear layer followed by a softmax function, the configuration of which transforms these embeddings into a probability distribution over the vocabulary, enabling the model to predict the next token in the sequence.

### v2

@@ 3.2.1 Implementation Details
Figure 3 illustrates the conceptual framework of our approach here. We concatenate the embeddings of two headline embeddings888 Because embeddings approximate the next token probability, one can also use the difference (subtraction) of two embedding vectors instead of the concatenation as the input to the downstream classification model. In our task, these two input types have no significantly different impact on the final classification accuracy. and apply the downstream classification models to predict the probability that the first headline will generate a higher CTR. We consider three embedding models:
  • $\bullet$
@@
    OpenAI embeddings: At the time of this study, the latest and best embedding model available from OpenAI was text-embedding-3-large. These embeddings are obtained using the OpenAI embedding API in Python. This embedding model generates embeddings with up to 3,072 dimensions and supports a maximum token limit of 8,191, with its knowledge base extending up to September 2021. A notable feature of this model is its flexibility – it allows users to tailor the embedding dimension to their specific needs. [^9] Other models capable of producing text embeddings, like BERT (devlin2018bert) and Sentence-BERT (reimers2019sentence), have fixed embedding vector dimensions that remain constant once the model is trained. These dimensions cannot be changed or resized for different applications. However, OpenAI’s embedding model uses a technique that, while producing outputs of fixed length, encodes information at different granularities (kusupati2022matryoshka). This allows a single embedding to adapt to the computational constraints of downstream tasks. More specifically, downstream tasks can use only the first several dimensions of the original vector, and the information retained decreases smoothly as a shorter part of the vector is used. In our experiments, we consider both 3,072- and 256-dimensional embeddings to examine whether the size of the embedding vector size affects the performance of the downstream classification models. [^10] A higher dimension does not necessarily mean better performance. While a higher dimension can retain more information for each sample, it also means that the samples are spread more sparsely in a higher-dimensional space given a limited number of training samples. This sparsity can increase the possibility of over-fitting.
  • $\bullet$
    Llama-3-8b: Meta’s Llama-3 (metallama3) is regarded as one of the most advanced open-source LLMs to date (llama3modelcard). We obtain embeddings from Llama-3 with 8 billion parameters (Llama-3-8b). We fine-tune this exact model using LoRA in §$\S$3.3; therefore, using the embeddings from this model allows for a direct comparison of fine-tuning vs. embedding-based classification methods (see §$\S$3.4 for comparisons). Llama-3-8b does not produce embeddings explicitly designed to represent the semantic meaning of entire input sequences. To address this issue, we apply the technique described in (behnamghader2024llm2vec), which transforms the Llama-3-8b model into a more suitable text embedding model, outputting embeddings with a dimension of 4,096.
  • $\bullet$
    Word2Vec embeddings: Word2Vec, proposed by mikolov_etal_2013a, was the leading embedding model before the widespread adoption of Transformers in the Natural Language Processing (NLP) literature. Unlike LLM-based text embeddings, Word2Vec does not use Transformers and cannot pay attention to surrounding token information. Embeddings from Word2Vec have been used in the recent marketing literature in a variety of use cases, including product attribute modeling (timoshenko_hauser_2019; wang_etal_2022), the effect of creativity on product popularity (sozuer_etal_2024). We compute the Word2Vec embedding vector for each word in the headline and average them to get the headline embedding, a common practice in Word2Vec. To ensure consistency with the dimensions used for OpenAI embeddings, we consider Word2Vec embeddings with 256 and 3,072 dimensions.
Next, to assess the impact of model complexity on performance, we consider two classification models:
  • $\bullet$
    Logit model with linear features: This is one of the simplest binary classification models and predicts the probability that a given input belongs to a particular class by applying the logistic function to a linear combination of the input variables.
  • $\bullet$
@@
    Multilayer perceptron: MLP is a powerful neural network architecture consisting of multiple linear and non-linear hidden layers that allow the model to learn complex patterns in the data. It employs a fully connected neural network with ReLU activation for hidden layers and Sigmoid transformation for the output probability, with dropout layers added to prevent over-fitting. We refer interested readers to Chapter 5 in (zhang2021dive) for more details. We train the MLP using the stochastic gradient descent algorithm to minimize binary cross-entropy and perform hyper-parameter tuning with Bayesian Optimization to determine the optimal configuration. Please see Web Appendix §$\S$§C.1 for additional details of the fine-tuning process and the architectures used in the final implementation.
Figure 3: The pipeline of the classification with the LLM text embeddings. Headline 1 and headline 2 are natural language sentences while Embedding 1 and Embedding 2 are numerical vectors.
Finally, in terms of the data, as discussed earlier, we choose all headline pairs that are significant at the 0.05 level. Note that the same headline can appear in multiple pairs when there are more than two headlines tested in one experiment. To avoid information leakage, we first randomly split the data into 80% training and 20% test sets based on the test ID (the experiment ID for Upworthy) to ensure all headline pairs under the same experiment are either in the training set or the test set. Notice that Upworthy allowed additional experiments for the same headline under a new experiment ID because the editor may require an extra test to make an informed decision. It means we may have the same headlines in both the training and test sets, even if they were in different experiments. In such cases, we discard headlines that have appeared in the training set from the test set to avoid information leakage. [^11] There are alternative ways to do the train-test split without discarding any samples. One approach is assigning one headline pair at one time to a training or test set and then adding all pairs with the same headline in the same group until all samples are allocated. However, after conducting several experiments in our study, we found that the splitting procedure has minimal impact on the final results. Hence, to maintain clarity and simplicity in our exposition, we simply discard around 25% of test headlines that have appeared in the training set.
@@ 3.2.2 Performance Analysis of Embedding-based Approaches
In Table 6, we present the performance of various embedding-based classification models.
@@
OpenAI-256E Logit 77.31%
MLP 77.62%
OpenAI-3072E Logit 82.50%
MLP 82.26%
Llama-3-8b-4096E Logit 74.03%
MLP 79.40%
Word2Vec-256E Logit 67.89%
MLP 67.32%
Word2Vec-3072E Logit 67.55%
MLP 67.02%
@@
Table 6: The performance of different embedding and classification models on test data. All models use the same training data, and accuracy is evaluated on the test data. Note that the Logit model’s training is a deterministic process after fixing the data, while MLP training incurs randomness via stochastic gradient descents. Therefore, we rerun MLP training three times with different sample paths (training set shuffling) and report the average accuracy from three reruns.
@@
At a high level, we see that all the embedding-based classification models perform better than prompt-based approaches. This performance can be attributed to the fact that these models are trained on a customized dataset, allowing them to learn decision rules tailored to the headline selection task. Because these models can leverage a larger labeled dataset during training, they are more generalizable and robust. In contrast, in-context learning relies on the LLM’s ability to understand and generalize from only a few demonstrations. Furthermore, the task of choosing the catchier headline from two options may not be a common scenario in OpenAI’s training corpus. Consequently, GPT models may struggle to generalize effectively from the limited examples provided in prompts, leading to less accurate predictions.
Next, we compare performance under different embedding models. We see that OpenAI embeddings outperform both Llama-3-8b and Word2Vec-based models. This is understandable since OpenAI’s LLMs are state-of-the-art models that naturally perform better than the classic Word2Vec embeddings. Also, the performance of embeddings from open-source Llama-3-8b falls between that of the proprietary OpenAI model and the older Word2Vec that does not use the transformer architecture. We also see that for OpenAI, higher-dimensional embeddings give higher-quality encoding and, thus, higher prediction accuracy. However, there is some limit beyond which there are diminishing returns. In our experiment, the 3072-dimensional embedding (OpenAI-3072E) improves over the 256-dimensional embedding (OpenAI-256E) by around 5%. A more striking observation is that classification model complexity does not improve the prediction for OpenAI embeddings. Note that there is no significant difference in accuracy between MLP and Logit for the following embedding models – OpenAI-256E, OpenAI-3072E, Word2Vec-256E, and Word2Vec-3072E. But for Llama-3-8b embeddings, the MLP classification model outperforms Logit classification model by approximately 5%. We hypothesize that with more dimensions in text embeddings, the separation between different classes becomes more complex (e.g., higher sparsity and higher risk of over-fitting), requiring a more sophisticated model for classification. Overall, our findings suggest that a Logit classifier with linear features on top of fixed text embeddings from OpenAI can effectively capture the attractiveness of content.
We now investigate a few other aspects of these models and summarize our findings below. For brevity, we simply discuss the findings here and refer interested readers to the Web Appendix for details. Note that in all these analyses, we focus on OpenAI embeddings (since it offers the best performance) and the logistic regression with linear features as our classification model (since the MLP does not offer any significant improvement over the logit model for OpenAI embeddings and is easier to train).
  • $\bullet$
    First, we examine the conditions under which these approaches fail. We use cosine similarity measures to quantify the similarity between the two headlines in a pair and find that these approaches tend to perform worse when the headlines are similar. This is understandable since text embeddings may not be able to capture the differences between two headlines when they are similar. As a result, it is hard for a classification model to differentiate between them. See Web Appendix §$\S$§C.2 for details.
  • $\bullet$
    Second, we quantify the role of the number of training samples on predictive accuracy. For the 256E-dimensional embedding, we find that approximately 65% of the training set, comprising approximately 25,000 pairs (with the remaining 35% used for the test set), is adequate to achieve high and stable performance. However, when we use the 3072E-dimensional embedding, there is evidence of overfitting even when we use 95% of the data for training. This suggests that for larger embedding vectors, we need much larger training datasets. See Web Appendix §$\S$§C.3 for details.
  • $\bullet$
    Third, we examine whether we can improve the performance and alleviate any overfitting issues by adopting dimensionality reduction techniques (PCA) and/or regularization (adding $L^{1}$/$L^{2}$ norms). However, we don’t see any significant improvement with these methods. See Web Appendix §$\S$§C.4 for details.
  • $\bullet$
    Finally, there are two potential threats to the inference from our analysis. First, one may be concerned as to whether Upworthy’s data has been used as part of OpenAI’s training corpus, which could potentially invalidate the results. A second concern is that readers’ preferences changed over time and that the current sample splitting method fails to take into account when the headline was created. We examine both these issues and provide a detailed discussion and robustness checks to alleviate concerns around them in Web Appendix §$\S$§C.5.
3.3 Fine-Tuning Open-Source LLMs with LoRA
@@
In §$\S$3.2, we trained classification models using text embeddings as fixed inputs, bypassing the need to train the LLM itself. The best-performing embedding-based model reaches an accuracy of 82.50% without modifying any parameters in the LLM. This raises a natural question: if we can train LLMs directly, can we achieve better performance? However, training LLMs from scratch is costly and resource-intensive. Therefore, in this section, we turn to fine-tuning techniques, which involve partially adjusting the parameters in the LLM to improve performance while keeping the computational cost manageable. Note that to fine-tune LLMs, we need to be able to access the underlying model parameters. This is only feasible for open-source LLMs, like BERT and Llama. [^12] Some proprietary LLMs offer black-box fine-tuning. For instance, OpenAI allows fine-tuning GPT-3.5, GPT-4/GPT-4o (with limited experimental access) through its API in a black-box manner. Users provide the training and test datasets and select the base GPT model, but are not provided any details of the fine-tuning process or the underlying model parameter. This approach also has a number of other drawbacks – the fine-tuning API is quite expensive compared to the open-source fine-tuning approaches, the fine-tuned model cannot be hosted on a local machine, data privacy concerns since the media firm’s proprietary data will need to be submitted to OpenAI. Given these drawbacks, it is preferable for firms to fine-tune open-source LLMs, as the process and data are fully under the control of the firm. However, for the completeness of comparison, we fine-tuned the ChatGPT-3.5 model on the training set using this API, with the default configuration provided by OpenAI, and the same training-test dataset split. We obtained an accuracy of only 79.80% on the test set. The goal of fine-tuning is to adapt a general-purpose LLM for specific downstream tasks through a process typically involving supervised learning using labeled datasets.
@@
For the fine-tuning task, we use Meta’s Llama-3-8b as our primary base model. zhao_etal_2024 consider a large-scale experiment where they fine-tune different open-source LLMs and show that for NLP tasks, including news headline generation, Llama-3 offers one of the best performance. Llama-3 offers two versions: one with 8 billion parameters (Llama-3-8b) and another with 70 billion parameters (Llama-3-70b). We opted for the smaller version since it is easier to fine-tune and, as we will see, even with this smaller model, we see evidence of over-fitting (see §$\S$3.3.3 for details). Hence, we avoid fine-tuning the 70-billion-parameter model, where over-fitting problems are likely to be exacerbated.
@@
The rest of this section is organized as follows. We present the LoRA method for efficient fine-tuning in §$\S$3.3.1. The goal is not to delve into every technical detail but to provide the necessary background information for a high-level understanding of fine-tuning. Then we detail our implementation in §$\S$3.3.2 and report our results in §$\S$3.3.3. Background information related to Transformer block, the fundamental component of most LLMs, including Llama-3, and the basic architecture of Llama-3 are given in Web Appendix §$\S$§D.1.